Grade 11Mechanics

Q.A projectile is fired horizontally from an inclined plane(of inclination 30 degree with horizontal)with speed 50m/s.If g=10m/s^2,the range measured along incline is
A)500m
B)1000/3m
C)200*root2m
D)100*root3m

jiten
8 Years agoGrade 11

1 Answer

Venkat
8 Years ago
There is some typo error on above solution. This may confuse you. Hence, I am reposting it .Answer is option (B)
T=\frac{2usin(\alpha+\beta)}{gcos\beta } \\ \\T=\frac{2usin\beta}{gcos\beta} \\ \\T=\frac{2*50*tan30}{10} = 57.735 \\ \\R=\frac{u^{2}*[sin(2\alpha+\beta)+sin\beta]}{gcos^{2}\beta} \\ \\R=\frac{u^{2}*[2sin\beta]}{gcos^{2}\beta} \\ \\R=\frac{uT}{cos\beta} \\ \\R=\frac{50*57.735}{cos30} \\ \\R=333.33 \\ \\R=\frac{1000}{3}.